OpenStudy (anonymous):

solve for x Cos x - sin x = 1

myininaya (myininaya):

well what is 1-0?

myininaya (myininaya):

or in other words when is cos(x)=1 and sin(x)=0

OpenStudy (anonymous):

at 0pi

myininaya (myininaya):

I know that would give one kinda solution we could also find a solution be considering 0-(-1)

myininaya (myininaya):

so 0+2pi*n where n is integer for when cos(x)=1 and sin(x)=0

myininaya (myininaya):

what answer would get considering 0-(-1) now

OpenStudy (anonymous):

1

OpenStudy (solomonzelman):

how about x=45?

myininaya (myininaya):

that isn't what i meant but yeah 0-(-1) is 1

OpenStudy (solomonzelman):

no nvm

myininaya (myininaya):

so we have cos(x)-sin(x)=1 and we know on the unit circle somewhere we will have 0 -(-1)=1 so when is cos(x)=0 and sin(x)=-1

OpenStudy (solomonzelman):

I thought =0, for some reason... don't know why. srry abt that

OpenStudy (anonymous):

sin is 0 at x= pi/2 and 3pi/2. cos is 0 at x= 0pi and pi

myininaya (myininaya):

sin is 1 at x=pi/2 and sin is -1 at x=3pi/2 sin is 0 at x=0,x=pi cos is 0 at x=pi/2 and 3pi/2 cos is 1 at x=0 cos is -1 at x=pi ----------- cos(x)-sin(x)=1 1-0=1 so cos(x)=1 when x=0,2pi n is integer cos(x)-sin(x)=1 0-(-1)=1 so sin(x)=-1 when x=3pi/2 if we are solving in the 0 to 2pi interval I don't see that pi is a solution though cos(pi)=-1

myininaya (myininaya):

you can skip the line where n is integer I was thinking about giving the general solution but I will let you come up with that

OpenStudy (anonymous):

thanks guys!! :D